# JailDAM: Jailbreak Detection with Adaptive Memory for Vision-Language Model

### v2

@@ 1 Introduction
GradSafe (Xie et al., 2024) Hidden states-based Zero
@@
2 Related Work
2.1 Jailbreak Detection and Defense on VLM
Recently, jailbreaking represents a critical concern for safety problem on language models (Chua et al., 2024; Yi et al., 2024; Xing et al., 2024; Yang et al., 2025). In the vision-language setting, several jailbreak detection methods have been proposed, which can be classified into four distinct categories: (1) LLM hidden states-based methods, (2) fine-tuning based methods, (3) prompt-based methods, and (4) uncertainty-based method. For jailbreak detection relying on hidden states, Huang et al. (2024) introduces a method called NEARSIDE, a detection method for VLMs that uses a single embedding vector (the ”attacking direction”) from the difference between harmful and benign dataset to efficiently classify harmful and benign inputs without requiring multiple inferences or expensive computations. Jiang et al. (2025) identifies safety-aware layers by analyzing hidden activations for safe and unsafe inputs, revealing layers that encode signals indicating whether a prompt is unsafe. It then constructs a Refusal Vector (RV) in the vocabulary space, capturing the model’s refusal behavior as a reference for detecting unsafe requests.
2.2 Test-time Adaptation
Test-time adaptation is the process of adjusting models to testing data that may differ in distribution from the training data. Various approaches have been proposed to address this challenge, which can be broadly categorized into methods that update model parameters, prompt-tuning techniques, and training-free adaptation strategies. One line of research focuses on updating a subset of model parameters to adapt to distribution shifts during testing. For instance, Yi et al. (2023) introduces TeCo, a test-time optimization framework that enhances video classification robustness by adjusting shallow-layer parameters while adapting batch normalization statistics in deeper layers. This allows the model to maintain stability while improving its generalization to unseen test data.
@@ 3.3 Memory Bank Generation
To build a robust safety mechanism, we construct a memory bank of unsafe concepts
@@
that serve as reference points for detecting harmful content(see Figure 2). The generation of these concepts is crucial for defining the boundaries of unsafe inputs. Given the vast and evolving nature of harmful content, we leverage GPT-4o (Achiam et al., 2023), a large-scale language model with broad knowledge of harmful content policies, to augment and generate structured representations of unsafe memories. Motivated by the idea from how people created interpretable concepts without human annotations in computer vision (Liu et al., 2024b; Oikarinen et al., 2023), we use GPT-4o one time to convert existing harmful content policies into a set of structured, high-level memories.(see Appendix B.1 for detailed generation prompt and concepts example) The process follows a structured prompt to extract n key concepts per jailbreak category, covering:
@@ 3.6 Application: An End-to-End Attack Defense Framework
Based on our attack detector, JailDAM . We also construct an end-to-end attack defense framework, denoted as JailDAM-D (see Figure 3). In this framework, we implement a two-stage defense approach. First, incoming instruction $S$ is analyzed by the JailDAM , which evaluates whether the input contains potential attack patterns or harmful content. If the JailDAM identifies the input as potentially harmful, the framework automatically concatenates a specialized defense prompt $T_{d}$ (see Appendix B.2) before the original query. This defense prompt serves to alert the target VLMs about the potential risks and instructs it to refuse answering the harmful request. Conversely, if the input is classified as benign, the query proceeds to the target VLMs without modification, allowing the target model to provide normal assistance.
@@ 4.1.1 Datasets
Our evaluation employs three established VLM jailbreak benchmarks: MM-SafetyBench (Liu et al., 2024a), FigStep (Gong et al., 2023), and JailBreakV-28K (Luo et al., 2024), covering diverse structure-based jailbreak samples. For each of the Jailbreak dataset, we pair it with the benign MM-Vet dataset (Yu et al., 2023). And we do the same for two tasks: (1) Jailbreak Detection and (2) Jailbreak Defense. Dataset details are provided in Appendix C.1 and C.2.
@@ 4.1.2 Baselines
For Jailbreak Detection, we evaluate JailDAM against diverse baselines across three categories: (1) uncertainty-based methods: JailGuard (Zhang et al., 2023a), (2) fine-tuning-based methods: Llavaguard (Helff et al., 2024) and VLGuard (Zong et al., 2024), and (3) hidden states-based methods: HiddenDetect (Jiang et al., 2025), GradSafeWe transfer and implement GradSafe (Xie et al., 2024), an LLM detection method, to VLM as our baseline approach. (Xie et al., 2024). For Jailbreak Defense, we compare JailDAM-D with existing prompt-based defense methods: Adashield (Wang et al., 2024) and FSD (Gong et al., 2023). Detailed detection prompts for fine-tuning-based methods can be found in Appendix D.1, and defense prompt settings are in Appendix D.2.
@@ 4.1.3 Metrics
For Jailbreak Detection, we use Area Under the Receiver Operating Characteristic curve (AUROC) and Area Under the Precision-Recall curve (AUPRC) to evaluate classifier performance across varying thresholds. For Jailbreak Defense, different from previous work that cares more on defense effectiveness (Wang et al., 2024), we design a confusion matrix (see
@@
Table 2) with jailbreak samples as the positive class and benign samples as the negative class. We then employ F1-score as primary metric. This metric provide a more comprehensive idea of how model balance between defense effectiveness and over-defensiveness given that over-defense is prevalent for lots of models (Li & Liu, 2024). Following Wang et al. (2024); Liu et al. (2024a) and Luo et al. (2024), we use GPT-4o-mini to detect refusal phrases (e.g., ”I’m sorry”, ”As an AI, I cannot”) in model responses using customized prompts tailored for jailbreak and benign inputs (see prompts in Appendix E). Responses containing refusal phrases are classified as ”Predict as Harmful”; otherwise, they are labeled as ”Predict as Benign”.
@@ 4.2 Main Results
Jailbreak Defense.
4.3 Analysis & Ablation Study
4.3.1 Time Cost Analysis
JailDAM excels with minimal training time.
4.3.2 Ablation on Concepts Size of Each Category
Figure 6: AUROC of detection task on different concept sizes
4.3.3 Ablation on Test-time Adaptation
Test-time adaptation significantly boosts JailDAM performance.
Based on our choice of concept size of 40 for each harmful content category (see§ 7), we conduct an ablation study on the
Adaptation Metric MS FS JBV
W/O AUROC 0.9044 0.8379 0.8526
AUPRC 0.8854 0.8119 0.7677
With AUROC 0.9472 ($\uparrow$) 0.9449 ($\uparrow$) 0.8734 ($\uparrow$)
AUPRC 0.9155 ($\uparrow$) 0.9121 ($\uparrow$) 0.8750 ($\uparrow$)
Table 5: Ablation study on adaptation effectiveness. In here, MS: MM-SafetyBench, FS: FigStep, JBV: JailBreakV-28K
5 Conclusion
6 Ethics Statement
7 Reproducibility Statement
@@
Appendix A Details of Related Work
A.1 Jailbreak Detection on VLM
A.2 Test-time Adaptation
Appendix B Supplementary of JailDAM and JailDAM-D
B.1 Concept Generation Prompt to Memory Bank
Figure 7: Concepts Generation Prompt
Figure 8: Concepts Samples
B.2 Defense Prompt of JailDAM-D
As shown in Figure 9, we design defense prompt for JailDAM-D pipeline:
Appendix C Details of Datasets
C.1 Details of Each Datasets
MM-SafetyBench.
FigStep.
JailbreakV-28k.
MM-Vet.
C.2 Dataset Statistic
Appendix D Details of Baseline
D.1 Jailbreak Detection
D.1.1 General Detection Prompt
Figure 11: Details of harmful content categories
Figure 12: General Detection Prompt
D.2 Jailbreak Defense
D.2.1 FSD
Figure 13: Defense Prompt of FSD
D.2.2 Adashield-S
Figure 14: Static Defense Prompt of Adashield-S (Wang et al., 2024)
Appendix E Details of Metrics
E.1 Judge Prompt for VLM response to Jailbreak data
Output: ”””
Figure 15: Attacking Check Prompt for GPT4o-mini
E.2 Judge Prompt for VLM response to Benign data
Output: ”””
Figure 16: Benign Refusing Check Prompt for GPT4o-mini

### v3

@@ 1 Introduction
GradSafe(vision) (Xie et al., 2024) Hidden states-based Zero
@@ 2.3 Memory Bank Generation
To build a robust safety mechanism, we construct a memory bank of unsafe concepts that serve as reference points for detecting harmful content(see Figure 2). The generation of these concepts is crucial for defining the boundaries of unsafe inputs. Given the vast and evolving nature of harmful content, we leverage GPT-4o (Achiam et al., 2023), a large-scale language model with broad knowledge of harmful content policies, to augment and generate structured representations of unsafe memories. Motivated by the idea from how people created interpretable concepts without human annotations in computer vision (Liu et al., 2024b; Oikarinen et al., 2023), we use GPT-4o one time to convert existing harmful content policies into a set of structured, high-level memories.(see Appendix C.1 for detailed generation prompt and concepts example) The process follows a structured prompt to extract n key concepts per jailbreak category, covering:
@@ 2.6 Application: An End-to-End Attack Defense Framework
Based on our attack detector, JailDAM . We also construct an end-to-end attack defense framework, denoted as JailDAM-D (see Figure 3). In this framework, we implement a two-stage defense approach. First, incoming instruction $S$ is analyzed by the JailDAM , which evaluates whether the input contains potential attack patterns or harmful content. If the JailDAM identifies the input as potentially harmful, the framework automatically concatenates a specialized defense prompt $T_{d}$ (see Appendix C.2) before the original query. This defense prompt serves to alert the target VLMs about the potential risks and instructs it to refuse answering the harmful request. Conversely, if the input is classified as benign, the query proceeds to the target VLMs without modification, allowing the target model to provide normal assistance.
@@ 3.1.1 Datasets
Our evaluation employs three established VLM jailbreak benchmarks: MM-SafetyBench (Liu et al., 2024a), FigStep (Gong et al., 2023), and JailBreakV-28K (Luo et al., 2024), covering diverse structure-based jailbreak samples. For each of the Jailbreak dataset, we pair it with the benign MM-Vet dataset (Yu et al., 2023). And we do the same for two tasks: (1) Jailbreak Detection and (2) Jailbreak Defense. Dataset details are provided in Appendix D.1 and D.2.
@@ 3.1.2 Baselines
For Jailbreak Detection, we evaluate JailDAM against diverse baselines across three categories: (1) uncertainty-based methods: JailGuard (Zhang et al., 2023a), (2) fine-tuning-based methods: Llavaguard (Helff et al., 2024) and VLGuard (Zong et al., 2024), and (3) hidden states-based methods: HiddenDetect (Jiang et al., 2025), GradSafeWe transfer and implement GradSafe (Xie et al., 2024), an LLM detection method, to VLM as our baseline approach. (Xie et al., 2024). For Jailbreak Defense, we compare JailDAM-D with existing prompt-based defense methods: Adashield (Wang et al., 2024) and FSD (Gong et al., 2023). Detailed detection prompts for fine-tuning-based methods can be found in Appendix E.1, and defense prompt settings are in Appendix E.2.
@@ 3.1.3 Metrics
We design a confusion matrix (see Table 2) with jailbreak samples as the positive class and benign samples as the negative class. For Jailbreak Detection, we use Area Under the Receiver Operating Characteristic curve (AUROC) and Area Under the Precision-Recall curve (AUPRC) to evaluate classifier performance across varying thresholds. For Jailbreak Defense, different from previous work that cares more on defense effectiveness (Wang et al., 2024), we employ F1-score as primary metric. This metric provide a more comprehensive idea of how model balance between defense effectiveness and over-defensiveness given that over-defense is prevalent for lots of models (Li & Liu, 2024). Following Wang et al. (2024); Liu et al. (2024a) and Luo et al. (2024), we use GPT-4o-mini to detect refusal phrases (e.g., ”I’m sorry”, ”As an AI, I cannot”) in model responses using customized prompts tailored for jailbreak and benign inputs (see prompts in Appendix F). Responses containing refusal phrases are classified as ”Predict as Harmful”; otherwise, they are labeled as ”Predict as Benign”.
@@ 3.2 Main Results
Jailbreak Defense.
3.3 Analysis & Ablation Study
3.3.1 Time Cost Analysis
JailDAM excels with minimal training time.
3.3.2 Ablation on Concepts Size of Each Category
Figure 6: AUROC of detection task on different concept sizes
3.3.3 Ablation on Memory Comprehensiveness
Memory comprehensiveness enhances cross-domain generalization.
The detector’s ability to generalize across different domains relies heavily on maintaining a comprehensive memory bank. Even when there is a domain gap between training and testing data, the model can effectively capture harmful content as long as the relevant harmful policies are well represented in the memory bank.
Adaptive Memory Mechanism ensures robustness under incomplete memory coverage.
To evaluate the resilience of our approach when memory coverage is incomplete, we perform a memory removal experiment on the MMSafetyBench dataset. Beginning with a full memory bank encompassing all 13 harmful categories, we progressively remove 1, 3, 5, and 7 categories and measure the corresponding detection performance. As reported in Table 4, both AUROC and AUPR metrics exhibit a gradual decline with increased removal of categories. The relatively slow degradation in performance indicates that our Adaptive Memory Mechanism effectively compensates for missing information, maintaining stable detection accuracy even when the memory bank is partially incomplete.
Metric Remove 0 Class Remove 1 Class Remove 3 Classes Remove 5 Classes Remove 7 Classes
AUROC 0.9472 0.9270 0.9140 0.9020 0.8350
AUPR 0.9155 0.9046 0.9068 0.8785 0.7917
Table 4: Ablation study on memory comprehensiveness on MM-SafetyBench.
3.3.4 Ablation on Test-time Adaptation
Test-time adaptation significantly boosts JailDAM performance.
Based on our choice of concept size of 40 for each harmful content category (see§ H), we conduct an ablation study on the
Adaptation Metric MS FS JBV
W/O AUROC 0.9044 0.8379 0.8526
AUPRC 0.8854 0.8119 0.7677
With AUROC 0.9472 ($\uparrow$) 0.9449 ($\uparrow$) 0.8734 ($\uparrow$)
AUPRC 0.9155 ($\uparrow$) 0.9121 ($\uparrow$) 0.8750 ($\uparrow$)
Table 5: Ablation study on adaptation effectiveness. In here, MS: MM-SafetyBench, FS: FigStep, JBV: JailBreakV-28K.
3.3.5 Ablation on Out of Distribution Benign Data
Benign Dataset Jailbreak Dataset AUROC AUPR
MM-Vet JailBreakV-28k 0.9465 0.9464
MMMU JailBreakV-28k 0.9034 0.8962
MM-Vet MM-SafetyBench 0.9472 0.9155
MMMU MM-SafetyBench 0.9452 0.9396
MM-Vet FigStep 0.9608 0.9616
MMMU FigStep 0.8852 0.8766
Table 6: Out-of-domain generalization of JailDAM evaluated by testing on MMMU benign inputs paired with various jailbreak datasets, with MM-Vet as baseline.
JailDAM demonstrates strong out-of-domain generalization.
To directly assess the detector’s ability to generalize to out-of-domain test inputs, we conduct an experiment using the MMMU (Yue et al., 2024) dataset as a domain distinct from the training dataset MM-vet. Unlike MM-vet, which focuses on general VQA data, MMMU consists primarily of academic-style content. We then apply the detector, trained solely on MM-vet, to the MMMU dataset to evaluate its ability to correctly classify benign inputs. The results, summarized in Table 6, demonstrate strong performance across three different jailbreak datasets used as harmful inputs, with AUROC and AUPR scores consistently high. This indicates that our model robustly captures generalizable, policy-aligned safety signals beyond the training distribution.
4 Conclusion
5 Ethics Statement
Appendix A Related Works
A.1 Jailbreak Detection on VLM
A.2 Test-time Adaptation
Appendix B Limitation and Future Work
This study focuses specifically on structure-based jailbreak attacks and does not encompass adversarial attacks such as perturbation-based or gradient-based methods. While this focused approach allows for in-depth analysis of structure-based vulnerabilities, it represents a limitation in the comprehensiveness of our detection framework. Future research will extend this work by developing more generalizable detection mechanisms capable of identifying both structure-based and adversarial attacks within a unified framework.
Appendix C Supplementary of JailDAM and JailDAM-D
C.1 Concept Generation Prompt to Memory Bank
Figure 7: Concepts Generation Prompt
Figure 8: Concepts Samples
C.2 Defense Prompt of JailDAM-D
As shown in Figure 9, we design defense prompt for JailDAM-D pipeline:
Appendix D Details of Datasets
D.1 Details of Each Datasets
MM-SafetyBench.
FigStep.
JailbreakV-28k.
MM-Vet.
D.2 Dataset Statistic
Appendix E Details of Baseline
E.1 Jailbreak Detection
E.1.1 General Detection Prompt
Figure 11: Details of harmful content categories
Figure 12: General Detection Prompt
E.2 Jailbreak Defense
E.2.1 FSD
Figure 13: Defense Prompt of FSD
E.2.2 Adashield-S
Figure 14: Static Defense Prompt of Adashield-S (Wang et al., 2024)
Appendix F Details of Metrics
F.1 Judge Prompt for VLM response to Jailbreak data
Output: ”””
Figure 15: Attacking Check Prompt for GPT4o-mini
F.2 Judge Prompt for VLM response to Benign data
Output: ”””
Figure 16: Benign Refusing Check Prompt for GPT4o-mini
Appendix G Additional Results
@@
Appendix H Reproducibility Statement